TOPOLOGICAL ARRANGEMENT OF CURVES OF DEGREE 6 ON CUBIC SURFACES IN ℝP^{3}

G. MIKHALKIN

Abstract. A quadric in ℝP^{3} cuts a curve of degree 6 on a cubic surface in ℝP^{3}. The papers classifies the nonsingular curves cut in this way on non-singular cubic surfaces up to homeomorphism.

Two issues new in the study related to the first part of the 16th Hilbert problem appear in this classification. One is the distribution of the components of the curve between the components of the non-connected cubic surface which turns out to depend on the patterns of arrangements (see Theorem 1). The other is presence of positive genus among the components of the complement and genusrelated restrictions (see Theorems 3 and 4).